# RWKV-CLIP: A Robust Vision-Language Representation Learner

### v1

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has significantly improved performance in various vision-language tasks by expanding the dataset with image-text pairs obtained from websites. This paper further explores CLIP from the perspectives of data and model architecture. To address the prevalence of noisy data and enhance the quality of large-scale image-text data crawled from the internet, we introduce a diverse description generation framework that can leverage Large Language Models (LLMs) to synthesize and refine content from web-based texts, synthetic captions, and detection tags. Furthermore, we propose RWKV-CLIP, the first RWKV-driven vision-language representation learning model that combines the effective parallel training of transformers with the efficient inference of RNNs. Comprehensive experiments across various model scales and pre-training datasets demonstrate that RWKV-CLIP is a robust and efficient vision-language representation learner, it achieves state-of-the-art performance in several downstream tasks, including linear probe, zero-shot classification, and zero-shot image-text retrieval. To facilitate future research, the code and pre-trained models are released at https://github.com/deepglint/RWKV-CLIP * Equal contribution. $\dagger$ Corresponding author.
@@ 1 Introduction
The proliferation of mobile networks and social platforms has greatly accelerated the large-scale production of image-text pairs. This unprecedented abundance of data has established the foundation for vision-language pre-training. Contrastive Language-Image Pre-training (CLIP) employs two distinct unimodal encoders for images and text, utilizing a contrastive loss, a highly effective mechanism for representation learning. Having been pre-trained on extensive image-text pairs collected from the internet, CLIP demonstrates strong transferability and has been widely applied across various domains Zhou et al. (2023); Yao et al. (2023).
@@ 2.1 Vision-Language Representation Learning
As the milestone in vision-language representation learning, CLIP Radford et al. (2021) has garnered unparalleled interest due to its remarkable zero-shot recognition capability and outstanding transfer performance. Subsequently, a significant amount of enhancement works based on CLIP have been proposed. SLIP Mu et al. (2022) combines self-supervised learning with CLIP pre-training to achieve significant performance improvements. DeCLIP Li et al. (2022b) employs multi-view supervision across modalities and nearest-neighbor supervision from similar pairs to enhance representation learning efficiency. FILIP Yao et al. (2022) refines contrastive loss to learn fine-grained representations for image patches and sentence words. UniCLIP Lee et al. (2022) boosts data efficiency by integrating contrastive loss across multiple domains into a single universal space. HiCLIP Geng et al. (2023) enhances cross-modal alignment by incorporating hierarchy-aware attention into both visual and language branches of CLIP. ALIP Yang et al. (2023) introduces a gating mechanism to reduce the influence of noisy pairs using synthetic data. Different from the above methods, this paper conducts further exploration of both the data and model architecture, proposing a diverse description generation framework and introducing RWKV-CLIP, the first RWKV-driven vision-language representation model.
2.2 Text Augumentation
@@ 3 Method
Figure 3: The architecture of RWKV-CLIP, which consists of M×$\times$ and N×$\times$ RKWV-driven blocks followed by an average pooling layer.
@@ 3.2 RWKV-CLIP
$\mathrm{aug}(T)=\mathrm{Sample}([T_{r},T_{s},T_{g}])$ (1)
@@
$\quad+\sum_{i=t+1}^{T-1}\text{diag}(\epsilon_{i,j})\cdot K_{s,i}^{\mathsf{T}}\cdot V_{s,i})$
@@
$\psi_{x}^{c}=\text{Lerp}_{\psi}(x)\cdot w_{\psi}^{c},\quad\psi\in\{R,K\}$ (8)
@@
$O_{x}^{c}=(\sigma(R_{x}^{c})\odot\rho(K_{x}^{c}))\cdot w_{o}^{c}$ (9)
@@
$\begin{split}L=-\sum_{i=1}^{N}\left[\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{i}^{\top}\hat{T}_{j}/\tau}}+\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{j}^{\top}\hat{T}_{i}/\tau}}\right]\end{split}$ (10)
@@ C.3 Case Study
a origami {label}. art of a {label}. graffiti of a {label}. a drawing of a {label}.
@@
a origami {label}. a low resolution photo of a {label}. the toy {label}. a rendition of the {label}.

### v2

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has significantly improved performance in various vision-language tasks by expanding the dataset with image-text pairs obtained from the web. This paper further explores CLIP from the perspectives of data and model architecture. To mitigate the impact of the noise data and enhance the quality of large-scale image-text data crawled from the internet, we introduce a diverse description generation framework that can leverage Large Language Models (LLMs) to combine and refine information from web-based image-text pairs, synthetic captions, and detection tags. Additionally, we propose RWKV-CLIP, the first RWKV-driven vision-language representation learning model that combines the effective parallel training of transformers with the efficient inference of RNNs. Extensive experiments across different model scales and pre-training datasets demonstrate that RWKV-CLIP is a robust vision-language representation learner and it achieves state-of-the-art performance across multiple downstream tasks, including linear probing, zero-shot classification, and zero-shot image-text retrieval. To facilitate future research, the code and pre-trained models are released at https://github.com/deepglint/RWKV-CLIP.
@@ 1 Introduction
The proliferation of mobile networks and social platforms has greatly accelerated the large-scale production of image-text pairs Yang et al. (2020); Yu et al. (2020). This unprecedented abundance of data has established the foundation for vision-language pre-training. Contrastive Language-Image Pre-training (CLIP) employs two distinct unimodal encoders for images and text, utilizing a contrastive loss, a highly effective mechanism for representation learning. Having been pre-trained on extensive image-text pairs collected from the internet, CLIP demonstrates strong transferability and has been widely applied across various domains Zhou et al. (2023); Yao et al. (2023).
@@ 2.1 Vision-Language Representation Learning
As the milestone in vision-language representation learning, CLIP Radford et al. (2021) has garnered unparalleled interest due to its remarkable zero-shot recognition capability and outstanding transfer performance. Subsequently, a significant amount of enhancement works Yang et al. (2024); An et al. (2024, 2023) based on CLIP have been proposed. SLIP Mu et al. (2022) combines self-supervised learning with CLIP pre-training to achieve significant performance improvements. DeCLIP Li et al. (2022b) employs multi-view supervision across modalities and nearest-neighbor supervision from similar pairs to enhance representation learning efficiency. FILIP Yao et al. (2022) refines contrastive loss to learn fine-grained representations for image patches and sentence words. UniCLIP Lee et al. (2022) boosts data efficiency by integrating contrastive loss across multiple domains into a single universal space. HiCLIP Geng et al. (2023) enhances cross-modal alignment by incorporating hierarchy-aware attention into both visual and language branches of CLIP. ALIP Yang et al. (2023) introduces a gating mechanism to reduce the influence of noisy pairs using synthetic data. Different from the above methods, this paper conducts further exploration of both the data and model architecture, proposing a diverse description generation framework and introducing RWKV-CLIP, the first RWKV-driven vision-language representation model.
2.2 Text Agumentation
@@ 3 Method
Figure 3: The architecture of RWKV-CLIP, which consists of M×$\times$ and N×$\times$ RWKV-driven blocks followed by an average pooling layer.
@@ 3.2 RWKV-CLIP
$\mathrm{aug}(T)=\mathrm{Sample}([T_{r},T_{s},T_{g}]).$ (1)
@@
$\quad+\sum_{i=t+1}^{T-1}\text{diag}(\epsilon_{i,j})\cdot K_{s,i}^{\mathsf{T}}\cdot V_{s,i}),$
@@
$\psi_{x}^{c}=\text{Lerp}_{\psi}(x)\cdot w_{\psi}^{c},\quad\psi\in\{R,K\}.$ (8)
@@
$O_{x}^{c}=(\sigma(R_{x}^{c})\odot\rho(K_{x}^{c}))\cdot w_{o}^{c},$ (9)
@@
$\begin{split}L=-\sum_{i=1}^{N}\left[\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{i}^{\top}\hat{T}_{j}/\tau}}+\log\frac{e^{\hat{I}_{i}^{\top}\hat{T}_{i}/\tau}}{\sum_{j}e^{\hat{I}_{j}^{\top}\hat{T}_{i}/\tau}}\right].\end{split}$ (10)
@@ 5 Conclusion
6 Limitations
Our proposed framework for diverse description generation leverages the existing caption generation model and detection tags model, both of which can directly influence the quality of the final generated descriptions. Furthermore, due to limitations in computational resources, this study only executes experiments at tens of millions of scales of image-text pairs. Conducting experiments at a billion-scale necessitates substantial computational resources.
@@ C.3 Case Study
an origami {label}. art of a {label}. graffiti of a {label}. a drawing of a {label}.
@@
an origami {label}. a low resolution photo of a {label}. the toy {label}. a rendition of the {label}.